OpenStudy (anonymous):

A metal forms two oxides . The lower oxide , M2O3 contains 30.4% oxygen . 7.9 g of lower oxides when converted to higher oxides become 10.3g. Ehich is the higher oxide?

OpenStudy (apoorvk):

Hmm. so, in M2O3, you have 30.4% O. so, the remaining 69.6% is the weight of the the two 'M' atoms right?

OpenStudy (anonymous):

yeah!

OpenStudy (anonymous):

ok this seems to be tought i will ask someone else plz look my next question!

OpenStudy (vincent-lyon.fr):

Let x be molar mass of M. \(\Large \frac{3\times 16}{2x+3\times 16}\normalsize =0.304\) solve for x